The question is: The perimeter of a rectangle is 96 meters. Three times the width is 8 meters less than the length. What are the dimensions of the rectangle?
A) Identify the two unknowns:
B) Write a system using the two unknowns:
C) Solve the system using any (Substitution, Elimination) method
I know that "Three times" is the same as saying 3x also that "the width is 8 meters less then the length" is the same as saying W=8-L. I know that P (Perimeter) is 96 meters. I just don't know how to solve for the rest of the problem!

Found 2 solutions by checkley77, edjones:
Answer by checkley77(12844)

L=3W-8
P=2L+2W
96=2(3W-8)+2W
96=6W-16+2W
96=8W-16
8W=96+16
8W=112
W=112/8
W=14 ANS. FOR THE WIDTH.
L=3*14-8
L=42-8
L=34 ANS. FOR THE LENGTH.
PROOF:
96=2*34+2*14
96=68+28
96=96

Answer by edjones(8007)

Comment: W=8-L is not the same as W=L-8
.
Let L=length, w=width and P=perimeter
2L+2w=P
2L+2w=96
.
3w=L-8
L=3w+8
.
2(3w+8)+2w=96 Substitution
6w+16+2w=96
8w=80
w=10 m
L=38 m
.
Ed